OpenStudy (anonymous):

what is the equation of a line that goes throught points (2,-1) and (4,-1). Please help thank you

OpenStudy (anonymous):

Given points \((x_1,y_1)\) and \((x_2,y_2)\), the equation of the line connecting them is given by the following formula.\[y-y_1=\frac{y_2-y_1}{x_2-x_1}(x-x_1)\]It doesn't matter which point gets the subscript 1 and the subscript 2, but it does matter that you are consistent through the entire application of the formula. This is a modification of the point slope equation of the line as the slope connecting the two points is given by \(m=\frac{\text{rise}}{\text{run}}=\frac{y_2-y_1}{x_2-x_1}\).

OpenStudy (anonymous):

thank you but where would i insert the y and x values?

OpenStudy (anonymous):

take the x and y of one point and replace the x_1 and y_1 take the x and y of the other point and replace the x_2 y_2 leave the x and y in the equation alone

OpenStudy (anonymous):

ok thank you guys